# Multiple Modes of Motion for the Effectiveness of Outer Hair Cells at High Frequencies

### v2

@@ Introduction
An example is shown in Fig. 10. Another examine is shown in Appendix B. The amplitude gain can exceed 100 fold for the LO and 80 fold for HO (Fig. 10A). These values are somewhat less in Fig. B1A. The amplitude gain generally increases with the operating point variable $\gamma$ as expected. This increase can be monotonic as in Fig. 10A. However, it can peak before $\gamma$ reaches unity as shown in Fig. B1A in Appendix B, depending on the set of parameter values.
The peak frequencies are similar for both oscillators (Fig. 10B). The LO is ahead of HO in phase, indicating the amplifying role of the LO even though the difference is rather small (Fig. 10C). The effect of the LO in affecting the amplitude of LO is quite large despite its much smaller mass and stiffness. These features are shared by the plots in Appendix B.
@@
Figure 10: An example of LO-driven elastic coupled oscillators (LOE). A: Amplitude gain over SO mode, B: Peak frequency, and C: The phase with respect to the external force, of each oscillator, is plotted against $\gamma$. LO (red) and HO (blue). D: The ratio of power output to power input. The set of parameter values: $c$=0.6, $\omega_{1}$=0.18 in Eq. 17.
@@
Figure 11: The amplitudes LO (red) and HO (blue) in LO-driven elastically coupled (LOE) oscillators. The abscissa is the frequency normalized to the resonance frequency $\omega/\omega_{r}$. The unit of the amplitude (ordinate axis) is $f/K$. The values of $\gamma$ are from 0.1(left) to 0.9(right) with an increment of 0.1.
@@ Discussion
The examination presented here is more like a survey rather than a thorough analysis due to the relatively large number of parameters. In addition, the focus is on the high frequency performance. Nonetheless it is possible to observe general features of the models.
@@
Amplitude gain of a coupled oscillator is in general an increasing function of $\gamma$, the the activity of the motile element in the OHC. Some of the cases (See Fig. 10A) show monotonic increase with $\gamma$ similar to the single mode oscillator with low elastic load (See Fig. 6A).
However, depending on the condition, it can peak before the operating point variable $\gamma$ reaches unity (See Figs. 8A and B1A in Appendix B).
@@
Short hair cells are more effective in the amplifying role with larger amplifying parameter $u$ because of their smaller $C_{0}$ and larger $k_{o}$. That is clearly illustrated by guinea pigs, which show a large length gradient from the base (20 $\mu$m) to the apex ($\sim 100\mu$m). However, mice show a much smaller gradient. The difference is only 1.5 times in length [39]. That indicates the significance of the structure of the organ of Corti in frequency tuning. Such a variability in cell length indicates the significance of the structural parameters and is thus compatible with the present analysis.
@@
A good number of macroscopic models of cochlear mechanics [4] assume that OHCs apply force directly on the basilar membrane (BM). Such models ignore issues resulting from microscopic structure, such as low-pass characteristics of intrinsic cellular electric circuit and impedance mismatch.
Paradoxically, the present analysis of coupled oscillator models shows a resemblance to the assumptions of those macroscopic models: The amplifier gain is not seriously constrained by the impedance mismatch. Both the amplitude and the phase of the two oscillators are not so large.
@@ Acknowledgments
I thank Drs. Richard Chadwick and Catherine Weisz of NIDCD for helpful comments and Drs. Matthew Kelley and Inna Belyantseva for reading the manuscript. This research was supported in part by the Intramural Research Program of the NIH, NIDCD.
@@ Appendix A Single oscillator with OHC
Here we consider a case, in which an OHC incorporated into a system with mechanical resonance as illustrated in Fig. A1. Described first is the case of electrical stimulation. Then the treatment is extended to mechanical stimulation. That is followed by a numerical examination similar to those examination of coupled oscillators that is presented in the main part.
@@
The description largely follows the one-dimensional model [13], rather than the more realistic cylindrical model [19] for simplicity.
@@
Let the motile element in the lateral membrane of the OHC have two states, the long state $L$ and the short state $S$, and let $P$ be the fraction of the motile element in state $L$. The natural length of the cell would be $X_{0}+aNP$, where $a$ is the contribution of a single motile element to the cell length accompanied by conformational change from $S$ to $L$. The same conformational change transfers a charge $q$ across the membrane. The number of the motile elements in the cell is $N$.
With an external elastic load $k_{e}$, the displacement $X$ due to the motile unit is given by
$X=aNP\;k_{e}/(k_{o}+k_{e}),$ (A1)
where $k_{o}$ is the structural stiffness of the cell. In addition, assume that $P_{\infty}$ at equilibrium follows the Boltzmann distribution
$P_{\infty}/(1-P_{\infty})=\exp[-\beta(G_{0}+qV+aF_{o})]$ (A2)
where $\beta=1/(k_{B}T)$ with Boltzmann’s constant $k_{B}$ and the temperature $T$. The constant term is represented by $G_{0}$. The quantities $V$ and $F_{o}$ are, respectively, the membrane potential and force applied from the external elastic load, which is associated with axial displacement $X$ of the cell. Since the stiffness of the external elastic load is e, $F_{o}=k_{e}X$.
@@
The equation of motion of the system that is illustrated in Fig. A1 can be expressed by
@@
where $X_{\infty}=aNP_{\infty}\;k_{o}/(k_{o}+K)$ is the displacement that corresponds to equilibrium, $m$ is the mass, and $\eta$ drag coefficient. The inertia term can be justified if the system is not far from equilibrium [19]. The difference between the current displacement $X$ and equilibrium displacement $X_{\infty}$ is the driving force. The equation of motion (A16) can be expresses using variable $P$
@@
$x=aNp\;k_{o}/(k_{o}+k_{e}),$ (A5)
in the manner exactly corresponding to $X$ and $P$, i.e. Eq. A1.
Figure A1: A simple system of OHC with mechanical load.
@@
where $e_{ec}$ is the endocochlear potential, $e_{K}$ is the resting potential of OHC, which is primarily determined by K+ conductance, and $R_{a}$ hair bundle conductance.
@@
From the transformed equation of motion Eq. A6 and equation for receptor potential Eq. A9, $p$ can be expressed as a linear function of $r$.
@@
Notice that Eqs. A15, A7, and A12 are, respectively, definitions of $u$, $u_{a}$, and $u_{q}$, which appear in the main text.
@@
The equation of motion in the presence of an external force $F$ is given by
@@
By following the same transformation as performed on voltage driven equation, this equation can be transformed into
@@
if the change in the hair bundle resistance $\hat{r}$ can be expressed $gx$ and $\hat{f}=f/(k+k_{o})$. This equation is exactly Eq. 8 in the main text.
Appendix B LOE another example
For LO-driven elastic coupled oscillators, peak gain increases with $\gamma$ while $\gamma$ is small. However, it may have a maximum before $\gamma$ reaches unity, depending on parameter sets. An example is shown below.
Figure B1: Another example of LO-driven elastic coupled oscillators (LOE). A: Amplitude gain over SO mode, B: Peak frequency, and C: The phase with respect to the external force, of each oscillator are plotted against $\gamma$. LO (red) and HO (blue). D: The ratio of power output to power input. The set of parameter values: $c$=0.4, $\omega_{1}$=0.24 in Eq. 17.

### v3

@@ Introduction
An example is shown in Fig. 10. The amplitude gain can exceed 100 fold for LO and 80 fold for HO (Fig. 10A). These values are somewhat less in Fig. 10A. The amplitude gain generally increases with the operating point variable $\gamma$ as expected. This increase can be monotonic as in Fig. 10A. However, it can peak before $\gamma$ reaches unity, similar to HO-driven case, depending on the set of parameter values.
The peak frequencies are similar for both oscillators (Fig. 10B). The LO is ahead of HO in phase, indicating the amplifying role of the LO even though the difference is rather small (Fig. 10C). The effect of the LO in affecting the amplitude of LO is quite large despite its much smaller mass and stiffness.
@@
Figure 10: An example of LO-driven elastically coupled oscillators (LOE). A: Amplitude gain over SO mode, B: Peak frequency, and C: The phase with respect to the external force, of each oscillator, is plotted against $\gamma$. LO (red) and HO (blue). D: The ratio of power output to power input. The set of parameter values: $c$=0.6, $\omega_{1}$=0.18 in Eq. 14.
@@
Figure 11: The amplitudes LO (red) and HO (blue) in LO-driven elastically coupled (LOE) oscillators. The peak values are plotted in Fig. 10A. The abscissa is the frequency normalized to the resonance frequency $\omega/\omega_{r}$. The unit of the amplitude (ordinate axis) is $f/K$. The values of $\gamma$ are from 0.1(left) to 0.9(right) with an increment of 0.1.
@@ Discussion
The numerical examination performed here concerns only high frequency performance. At lower frequencies, such as 10 kHz, viscosity coupled models perform much better than at high frequencies (not shown).
@@
Amplitude gain of a coupled oscillator is in general an increasing function of $\gamma$, the the activity of the motile element in the OHC. Some of the cases (See Fig. 10A) show monotonic increase with $\gamma$ similar to the single mode oscillator with low elastic load (See Fig. 6A). However, depending on the condition, it can peak before the operating point variable $\gamma$ reaches unity (See Figs. 8A).
@@
Short hair cells are more effective in the amplifying role having larger values of amplifying parameter $u$ because of their smaller $C_{0}$ and larger $k_{o}$. For this reason, basal cells, which operate high frequencies are short. For example, basal cells in both guinea pigs and mice are short ($\sim 20\;\mu$m long), even though their length gradient is steeper for guinea pigs [45].
@@
A good number of macroscopic models of cochlear mechanics [4] assume that OHCs apply force directly on the BM. Such models ignore issues resulting from microscopic structure, such as low-pass characteristics of intrinsic cellular electric circuit and impedance mismatch.
Interestingly, the present analysis of coupled oscillator models shows a resemblance to the assumptions of those macroscopic models: The amplifier gain is not seriously constrained by the impedance mismatch. Both the amplitude and the phase of the two oscillators are not so large.
@@ Conclusions
Author Contributions
The single author performed all, including conceptualization, development of the equations, computation using Mathematica, and composing the paper.
@@ Acknowledgments
I thank Drs.  Richard Chadwick and Catherine Weisz of NIDCD for helpful comments. Dr. Chadwick pointed out the significance of Eq. 28 of Wang et al [23]. I also appreciated help by Drs. Matthew Kelley and Inna Belyantseva for reading the manuscript. This research was supported in part by the Intramural Research Program of the NIH, NIDCD.
Declaration of interests
The author declares no competing interests.
@@ Appendix A Single oscillator with OHC
Here we consider a case, in which an OHC incorporated into a system with mechanical resonance as illustrated in Fig. A1. Described first is the case of electrical stimulation. Then the treatment is extended to mechanical stimulation.
@@
The description largely follows the one-dimensional model for electrical stimulation [14], rather than the more realistic cylindrical model [25] for simplicity because the descriptions of OHCs by these two models are consistent.
@@
Assume that the motile element in the lateral membrane of the OHC have two states, the long state $L$ and the short state $S$, and $P$ is the fraction of the motile element in state $L$. The natural length of the cell is $X_{0}+aNP$, where $a$ is the contribution of a single motile element to the cell length associated with conformational change from $S$ to $L$. This conformational change accompanies movement of a charge $q$ across the membrane. The number of the motile elements in the cell is $N$. In the equilibrium condition, $P$ is determined by the free energy $G$, which is the sum of an electrical term $qV$ and a mechanical term $aF_{e}$, where $V$ is the membrane potential (or voltage) and $F_{e}$ the force applied to the cell due to mechanoelectric coupling [57, 58].
Now put mechanical load, including an external elastic element $k_{e}$, a viscous element $\eta$, and a mass $m$, on an OHC with intrinsic stiffness $k_{o}$ (Fig. A1). Assume that this system is initially in equilibrium, and then change the membrane potential by $\Delta V$. The motile element undergoes conformational change $\Delta P$. The electrical component of the free energy change is $q\Delta V$. The resulting mechanical displacement $\Delta X_{p}=aN\Delta P$ generates force $\Delta F_{e}$ due to the external elastic load. This force, in turn, produces an elastic displacement $\Delta X_{e}$ in the OHC, resulting in the net displacement $\Delta X=aN\Delta P-\Delta X_{e}$ on the OHC as well as the external elastic element $K$. Thus, $\Delta F_{e}=k_{e}\Delta X=k_{o}\Delta X_{e}$. These relationships leads to $\Delta X=aN\Delta P\;k_{o}/(k_{o}+k_{e})$ and
$\Delta F_{e}=aN\Delta P\;k_{e}k_{o}/(k_{o}+k_{e}).$ (A1)
Thus, the total free energy is $q\Delta V+a\Delta F_{e}$ with $\Delta F_{e}$ given by Eq. A1.
We proceed by assuming that the free energy of the motile element at any given moment is determined by the given values of voltage $V$ and the mechanical strain $X$ of that moment, exactly the same as in the static case as described above. Then the variable $P$ of the motile element changes toward its equilibrium value $P_{\infty}$, which is given by the Boltzmann function
$P_{\infty}=\exp[-\beta\Delta G]/(1+\exp[-\beta\Delta G]),$ (A2)
with
$\Delta G=G_{0}+q\Delta V+aN\Delta P\;k_{e}k_{o}/(k_{o}+k_{e}),$ (A3)
and $\beta=1/(k_{B}T)$ with Boltzmann’s constant $k_{B}$ and the temperature $T$, and the constant term is represented by $G_{0}$. The difference $P-P_{\infty}$ drives the system.
@@
The equation of motion of the system can be formally expressed by
@@
where $X_{\infty}=aNP_{\infty}\;k_{o}/(k_{o}+K)$ is the displacement that corresponds to equilibrium. Here $m$ is the mass, $\eta$ drag coefficient, and $F$ external force. The inertia term can be justified if the system is not far from equilibrium [25]. This equation can be expresses using variable $P$
@@
$x=aNp\;k_{o}/(k_{o}+k_{e}).$ (A6)
Figure A1: Two simple oscillators, A and B, each of which consists of an OHC with intrinsic stiffness $k_{o}$, elastic load $k_{e}$ and viscous load $\eta$. The OHC responds to the movement of mass (dashed red arrow). The equations developed for A [24, 25] are equally applicable to B for electrical stimulation.
@@
where $e_{ec}$ is the endocochlear potential, $e_{K}$ is the resting potential of OHC, which is primarily determined by K+ conductance, and $R_{a}$ hair bundle conductance. The last term in the right-hand-side of the equation is due to the change of the motile mechanism.
@@
From the transformed equation of motion Eq. A7 and equation for receptor potential Eq. A10, $p$ can be expressed as a linear function of $\hat{r}$.
@@
Notice that Eqs. A16, A8, and A13 are, respectively, definitions of $u$, $u_{a}$, and $u_{q}$, which appear in the main text.
@@
The equation of motion in the presence of an external force $F$ can be formally given by
@@
If the motile mechanism in OHCs is as sensitive to applied force as to voltage changes, force stimulation requires a treatment that is symmetric to voltage stimulation. However, OHCs are a semi-piezoelectric resonator rather than a full piezoelectric resonator because he indirect effect of displacement via hair bundle conductance is more than an oder of magnitude greater than the direct effect of applied displacement on OHCs [39].
For this reason, the equation of motion can be expressed simply adding the external force as an additional term to the equation of motion developed for voltage stimulation. This applies to both models in Fig. A1. We may expect Model B responds to an external force similar to Model A under the condition $k_{e}\lesssim k_{o}$, but such a condition is not needed because of the high sensitivity of the hair bundle.
Thus, we arrive the equation in the following form in both cases:
@@
if the change in the hair bundle resistance $\hat{r}$ can be expressed $gx$ and $\hat{f}=f/(k_{e}+k_{o})$. This equation is Eq. 5 in the main text.